# Efficient quantum Gibbs samplers with Kubo–Martin–Schwinger detailed balance condition

### v2

@@ 3. A family of efficient quantum Gibbs samplers
In this section, we present a general framework for designing efficient quantum Gibbs samplers via Lindblad dynamics satisfying $\sigma_{\beta}$-KMS DBC.
@@ 3.1. Quantum Gibbs samplers via KMS-detailed balanced Lindbladian
Thanks to 10, the class of KMS detailed balanced Lindbladians can be parameterized by
@@
Note that the condition (2.28) is equivalent to $\Delta_{\sigma_{\beta}}^{-1/4}L_{j}=\Delta_{\sigma_{\beta}}^{1/4}L_{j}^{\dagger}$, namely, $\Delta_{\sigma_{\beta}}^{-1/4}L_{j}$ is self-adjoint. Thus, the admissible set of jump operators is
@@
Suppose that we are given a set of self-adjoint coupling operators $\{A^{a}\}_{a\in\mathcal{A}}$. For each $a$, we choose a weighting function $q^{a}(\nu):\mathbb{R}\rightarrow\mathbb{C}$, which satisfies:
@@
Here $\widehat{f^{a}}(\nu)$ can be viewed as a filtering function in the frequency domain, and its Fourier transform gives the filtering function in the time domain:
(3.3) $f^{a}(t)=\frac{1}{2\pi}\int^{\infty}_{-\infty}q^{a}(\nu)e^{-\beta\nu/4}e^{-it\nu}\,\mathrm{d}\nu\,.$
The choice of $q^{a}$ is a key component in our algorithm and will be discussed in detail in Section 3.2. Moreover, for each pair $(A^{a},q^{a})$, we define an operator $\widetilde{A}^{a}:=\sum_{\nu\in B_{H}}\widetilde{A}^{a}_{\nu}$ with $\widetilde{A}^{a}_{\nu}=q^{a}(\nu)A^{a}_{\nu}$, which can be easily verified to be self-adjoint:
$(\widetilde{A}^{a})^{{\dagger}}=\sum_{\nu\in B_{H}}(\widetilde{A}^{a})^{{\dagger}}_{\nu}=\sum_{\nu\in B_{H}}\overline{q^{a}(\nu)}(A^{a}_{\nu})^{{\dagger}}=\sum_{\nu\in B_{H}}q^{a}(-\nu)A^{a}_{-\nu}=\widetilde{A}^{a}\,.$
Then the jump operator for the Lindbladian, defined by
(3.4) $L_{a}=\sum_{\nu\in B_{H}}q^{a}(\nu)e^{-\beta\nu/4}A^{a}_{\nu}=\sum_{\nu\in B_{H}}\int^{\infty}_{-\infty}f^{a}(t)A^{a}_{\nu}e^{i\nu t}\,\mathrm{d}t=\int^{\infty}_{-\infty}f^{a}(t)A^{a}(t)\mathrm{d}t\,,$
satisfies the requirement in (2.28). We see that each $L_{a}$ is a linear combination of the Heisenberg evolution $A^{a}(t)=e^{iHt}A^{a}e^{-iHt}$.
Remark 12.
Noting that $L_{a}$ in Eq. 3.4 is the same as $\widehat{A}^{a}_{f^{a}}(\omega=0)$ in Eq. 1.2, if we let $f^{a}=f$ for all $a$, then the dissipative part of Lindbladian constructed via $\{L_{a}\}_{a\in\mathcal{A}}$ fits into the ansatz (1.2) with $\gamma(\omega)=\delta(\omega)$. This reduces a continuously parameterized set of jump operators to a discrete set.
@@
$=\sum_{a\in\mathcal{A}}\sum_{\nu,\nu^{\prime}\in B_{H}}\widehat{g^{a}}(\nu,\nu^{\prime})\left(A^{a}_{\nu^{\prime}}\right)^{\dagger}A^{a}_{\nu}\,,$
with the coefficient function $\widehat{g^{a}}$ in the frequency domain given by
(3.5) $\widehat{g^{a}}(\nu,\nu^{\prime})=\frac{1}{2i}\tanh(\beta(\nu^{\prime}-\nu)/4)e^{-\beta(\nu+\nu^{\prime})/4}q^{a}(\nu)\overline{q^{a}(\nu^{\prime})}\,.$
We define $g^{a}$ in the time domain via a two-sided Fourier transform as follows:
(3.6) $g^{a}(t,t^{\prime})=\frac{1}{(2\pi)^{2}}\iint_{\mathbb{R}^{2}}\widehat{g^{a}}(\nu,\nu^{\prime})e^{-i(\nu t-\nu^{\prime}t^{\prime})}\,\mathrm{d}\nu\,\mathrm{d}\nu^{\prime}\,.$
Since $\left\lvert\tanh(\cdot)\right\rvert\leq 1$ and $q^{a}(\nu)e^{-\beta\nu/4}\in L^{1}(\mathbb{R})$, $g^{a}(t,t^{\prime})$ is well defined. Then, it is direct to compute, by $(A^{a})^{\dagger}=A^{a}$ and (2.14),
(3.7) $\begin{split}G=&\sum_{a\in\mathcal{A}}\sum_{\nu,\nu^{\prime}\in B_{H}}\widehat{g^{a}}(\nu,\nu^{\prime})\left(A^{a}_{\nu^{\prime}}\right)^{\dagger}A^{a}_{\nu}\\ =&\sum_{a\in\mathcal{A}}\sum_{\nu,\nu^{\prime}\in B_{H}}\iint_{\mathbb{R}^{2}}e^{i\nu t}g^{a}(t,t^{\prime})e^{-i\nu^{\prime}t^{\prime}}A^{a}_{-\nu^{\prime}}A^{a}_{\nu}\,\mathrm{d}t\,\mathrm{d}t^{\prime}\\ =&\sum_{a\in\mathcal{A}}\iint_{\mathbb{R}^{2}}g^{a}(t,t^{\prime})A^{a}(t)A^{a}(t^{\prime})\,\mathrm{d}t\,\mathrm{d}t^{\prime}\,.\end{split}$
Therefore, letting $G$ and $\{L_{a}\}_{a\in\mathcal{A}}$ be constructed in (3.7) and (3.4), respectively, the Lindbladian
@@
Remark 13.
To ensure that the constructed Lindblad dynamics (3.8) eventually relaxes to the desired Gibbs state, by 3, we should carefully choose the coupling operators $A^{a}$ and weighting functions $q^{a}$ such that the resulting $\{L^{a}\}$ and $G$ satisfy Eq. 2.5. This is always possible, due to the finite dimensionality of the system. Moreover, it is known [TKR+10] that for a primitive KMS detailed balanced QMS, the mixing time (1.3) can be characterized by the spectral gap of the Lindbladian. The recent work [RFA24] estimated the spectral gap of the efficient quantum Gibbs sampler in [CKG23] in the high-temperature regime, by mapping the Lindbladian $\mathcal{L}_{\beta}$ with KMS DBC to a Hamiltonian ~$\widetilde{\mathcal{L}}_{\beta}:=\sigma_{\beta}^{-1/4}\mathcal{L}_{\beta}(\sigma_{\beta}^{1/4}X\sigma_{\beta}^{1/4})\sigma_{\beta}^{-1/4}$ and then analyzing its spectral properties with perturbation theory. The extension of such a mixing time analysis framework to our case with the optimal selection of $\{(A^{a},q^{a})\}_{a\in\mathcal{A}}$ will be discussed in a forthcoming work.
We have discussed the choice of self-adjoint coupling operators $\{A^{a}\}_{a\in\mathcal{A}}$ above. In fact, one can also generally consider a set of couplings such that $\{A^{a}\}_{a\in\mathcal{A}}=\{(A^{a})^{\dagger}\}_{a\in\mathcal{A}}$, and construct the corresponding jump operators $\{L_{a}\}_{a\in\mathcal{A}}$ and the Lindbladian $\mathcal{L}$ as in Eqs. 3.4 and 3.8, respectively. It is easy to see that the Lindblad dynamics defined in this way still satisfies the KMS detailed balance. Indeed, let $L_{a}$ and $L_{a,{\rm adj}}$ be the jumps associated with some $A^{a}$ and $(A^{a})^{\dagger}$ by (3.4) (without loss of generality, $A^{a}\neq(A^{a})^{\dagger}$). We then define self-adjoint operators
@@
3.2. Choice of the weighting function $q(\nu)$
In order to efficiently implement the jump operators $\{L_{a}\}_{a\in\mathcal{A}}$ in (3.4) and the coherent term $G$ in (3.7), we need to approximate the involved integrals on $\mathbb{R}$ and $\mathbb{R}^{2}$ by a numerical quadrature in a finite region. This requires $f^{a}(t),g^{a}(t,t^{\prime})$ to be smooth functions that decay rapidly as $\left\lvert t\right\rvert,\left\lvert t^{\prime}\right\rvert\to\infty$. To this end, we assume that $q^{a}$ is a compactly supported Gevrey function. We first recall the definition of Gevrey functions below [AHR17].
@@ 3.2. Choice of the weighting function $q(\nu)$
Some useful properties of Gevrey functions are collected in Appendix C. In particular, the product of two Gevrey functions is a Gevrey function (26); Certain compositions of Gevrey functions are Gevrey functions (28). The Fourier transform of compactly supported Gevrey functions satisfies Paley-Wiener type estimates (29).
@@
For $\beta>0$, suppose that $q(\nu)$ is a weighting function of the form $q(\nu)=u(\beta\nu)w(\nu)$ with the following conditions:
@@
  • (Compact support) There exists $S>0$ such that $\mathrm{supp}(w)\subset[-S,S]$.
  • (Gevrey) There exists $A_{q},A_{u},A_{w}\geq 1$, $s_{u},s_{w}\geq 1$ such that
    $u(\nu)e^{-\nu/4}\in\mathcal{G}^{s_{u}}_{A_{q},A_{u}}(\mathbb{R})\,,\quad w\in\mathcal{G}^{s_{w}}_{A_{q},A_{w}}(\mathbb{R})\,.$
@@
For the weighting function in 15, there holds (26)
@@
(3.12) $q(\nu)e^{-\beta\nu/4}\in\mathcal{G}^{s,}_{A_{q}^{2},\beta A_{u}+A_{w}}(\mathbb{R})\,,\quad s:=\max\{s_{u},s_{w}\}\,.$
Intuitively, one may expect that the functions $u(\nu)$ and $w(\nu)$ control the magnitude and support of the energy transition induced by a jump operator $L$, respectively. We can prove that the associated filtering functions $f(t)$ and $g(t,t^{\prime})$ in the time domain decay rapidly (30). This further allows us to show that a simple quadrature scheme (trapezoidal rule) can efficiently approximate $\{L_{a}\}_{a\in\mathcal{A}}$ and $G$ with high accuracy. Specifically, given $M=2^{\mathfrak{m}-1}$ with $\mathfrak{m}\in\mathbb{N}_{+}$ and $\tau>0$, the quadrature points are given by
@@
The quadrature error can be controlled as follows, and its proof is given in Appendix D.
@@
Under 15, we assume $\beta>0$, $\|A^{a}\|\leq 1$ for any $a\in\mathcal{A}$. When $\tau<\frac{2\pi}{2\left\lVert H\right\rVert+S}$ and
(3.14) $(M-1)\tau=\Omega\left((\beta A_{u}+A_{w})\log\left(\beta A_{u}+A_{w}\right)\right)\,,$
Then, it holds that
(3.15) $\left\|L_{a}-\sum^{2M-1}_{m=0}f^{a}(t_{m})A^{a}(t_{m})\tau\right\|\leq C_{f}\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2(\beta A_{u}+A_{w})^{1/s}e}\right)\,,$
@@
(3.16) $C_{f}=\mathcal{O}\left(A_{q}^{2}S(\beta A_{u}+A_{w})^{1/s}\right)\,,$
@@
(3.17) $\left\|G-\sum_{a\in\mathcal{A}}\sum^{2M-1}_{n,m=0}g^{a}(t_{n},t_{m})A^{a}(t_{n})A^{a}(t_{m})\tau^{2}\right\|\leq C_{g}|\mathcal{A}|\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2(2\beta A_{u}+2A_{w}+\beta)^{1/s}e}\right)\,,$
@@
(3.18) $C_{g}=\mathcal{O}\left(A_{q}^{4}S^{2}(\beta A_{u}+A_{w})^{1/s}\right)\,.$
Thanks to 16, to approximately block encode $\{L_{a}\}$ and $G$, it suffices to construct block-encodings for the discretized quantities
@@
$\sum_{a\in\mathcal{A}}\sum^{2M-1}_{n,m=0}g^{a}(t_{n},t_{m})e^{iHt_{n}}A^{a}e^{-iH(t_{n}-t_{m})}A^{a}e^{-iHt_{m}}\tau^{2}\,.$
In our algorithm, we construct these two block encodings using LCU (see Appendix B). This utilizes block encodings of $A^{a}$ (3.23), controlled Hamiltonian simulation (3.24), and prepare oracles for $f$ (Eqs. 3.26 and 3.27) or $g$ (Eqs. 3.29 and 3.30). The detailed constructions are presented in the next subsection (see (3.32) and (3.33)).
@@
Bounding the approximation error for $\{L_{a}\}_{a\in\mathcal{A}}$ and $G$ in the operator norm is a nontrivial task. [CKBG23] introduces a “rounding Hamiltonian” technique to bound the quadrature error in the frequency domain. By choosing weighting functions in 15, we can use the Poisson summation formula to simplify the quadrature error analysis. In particular, we can bound the quadrature error in the time domain without using the “rounding Hamiltonian” technique.
According to [AHR17, Corollary 2.8], for any $s_{w}>1$, there exists a “bump function” $w\in\mathcal{G}^{s_{w}}_{C_{w},A_{w}}$ such that $\mathrm{supp}(w)\subset[-1,1]$ and $w(\nu)=1$ when $|\nu|\leq 1/2$. Here $S$ is an adjustable parameter to control the support of $q$. Then $w(\nu/S)$ is supported on $[-S,S]$ and $w(\nu/S)\in\mathcal{G}^{s_{w}}_{C_{w},A_{w}/S}$. Now we provide two specific examples of $q$ satisfying 15.
@@
    where $u(\nu)e^{-\frac{\nu}{4}}=e^{-\frac{\sqrt{1+\nu^{2}}+\nu}{4}}\in\mathcal{G}^{1}_{1,7/2}$ is a Gevrey function of order $s_{u}=1$ and its derivative is $L^{1}$-integrable; see 28. Therefore $s=\max\{s_{u},s_{w}\}=s_{w}$. When $\beta\gg 1$, $q(\nu)$ in Eq. 3.19 gives a smoothed version of the Metropolis-type filter (similar to the Glauber-type filter):

### v4

@@ 3. A family of efficient quantum Gibbs samplers
In this section, we present a general framework for designing efficiently simulable quantum Gibbs samplers via Lindblad dynamics satisfying $\sigma_{\beta}$-KMS DBC, based on the representation 10.
@@ 3.1. Quantum Gibbs samplers via KMS-detailed balanced Lindbladian
By 10, the class of KMS detailed balanced Lindbladians can be parameterized by
@@
We first reformulate the condition (2.28) into a more convenient form. Note that (2.28) is equivalent to $\Delta_{\sigma_{\beta}}^{-1/4}L_{j}=\Delta_{\sigma_{\beta}}^{1/4}L_{j}^{\dagger}$, namely, $\Delta_{\sigma_{\beta}}^{-1/4}L_{j}$ is self-adjoint. Then the admissible set of jump operators can be given by
@@
Suppose that we are given a set of self-adjoint coupling operators $\{A^{a}\}_{a\in\mathcal{A}}$ as predefined jump proposals (coupling operators), where $\mathcal{A}$ is a discrete index set. For each $a\in\mathcal{A}$, we choose a weighting function $q^{a}(\nu):\mathbb{R}\rightarrow\mathbb{C}$ such that
@@
Here $q^{a}$ is chosen to be a symmetric function so that $L_{a}$ is an admissible jump (i.e., satisfies the condition (2.28)), see Eq. 3.5.
We then construct the jump operator associated with the pair $(A^{a},q^{a})$:
(3.3) $L_{a}:=\sum_{\nu\in B_{H}}q^{a}(\nu)e^{-\beta\nu/4}A^{a}_{\nu}=\sum_{\nu\in B_{H}}\widehat{f^{a}}(\nu)A^{a}_{\nu}=\int^{\infty}_{-\infty}f^{a}(t)A^{a}(t)\mathrm{d}t\,,$
as a weighted integral of $A^{a}$ (see Eq. 2.15), where the Fourier transform of $\widehat{f^{a}}$:
(3.4) $f^{a}(t)=\frac{1}{2\pi}\int^{\infty}_{-\infty}q^{a}(\nu)e^{-\beta\nu/4}e^{-it\nu}\,\mathrm{d}\nu$
gives the filtering function in the time domain. To verify that $L_{a}$ satisfies the condition (2.28), it suffices to note that $\widetilde{A}^{a}_{\nu}:=q^{a}(\nu)A^{a}_{\nu}$ corresponds to a self-adjoint operator defined by $\widetilde{A}^{a}:=\sum_{\nu\in B_{H}}\widetilde{A}^{a}_{\nu}$:
(3.5) $(\widetilde{A}^{a})^{{\dagger}}=\sum_{\nu\in B_{H}}(\widetilde{A}^{a})^{{\dagger}}_{\nu}=\sum_{\nu\in B_{H}}\overline{q^{a}(\nu)}(A^{a}_{\nu})^{{\dagger}}=\sum_{\nu\in B_{H}}q^{a}(-\nu)A^{a}_{-\nu}=\widetilde{A}^{a}\,.$
We see from (3.3) that each $L_{a}$ is a linear combination of the Heisenberg evolution $A^{a}(t)=e^{iHt}A^{a}e^{-iHt}$, and by (2.16), $L_{a}$ has an alternative representation:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L_{a}=\widehat{f^{a}}(-\beta^{-1}\log\Delta_{\sigma_{\beta}})A^{a}\,.}$
Here, the filtering function $\widehat{f^{a}}(\nu)$ is used to modify the transition between eigenvectors of $H$ with an energy difference of $\nu$, similar to the rejection step in the classical sampling, and the associated operator $\widehat{f^{a}}(-\beta^{-1}\log\Delta_{\sigma_{\beta}})$ maps the predefined coupling operators $A^{a}$ to the desired jumps $L_{a}$ for the $\sigma_{\beta}$-KMS detailed balance. Moreover, due to the introduction of $q^{a}$ in $L_{a}$, without loss of generality, we can let $\left\lVert A^{a}\right\rVert\leq 1$ by rescaling the norm.
@@
$=-\frac{i}{2}\sum_{a\in\mathcal{A}}\sum_{\nu\in B_{H}}\tanh\left(-\frac{\beta\nu}{4}\right)\left(L_{a}^{\dagger}L_{a}\right)_{\nu}\,,$
where $\left(L_{a}^{\dagger}L_{a}\right)_{\nu}$ is defined as in (2.13). However, the function $\tanh\left(-\beta\nu/4\right)$ is not $L^{1}$-integrable, and hence the useful time integral formulation as the one (3.3) for the jump cannot readily follow. To remedy this issue, thanks to $|\nu|\leq 2\|H\|$ for a Bohr frequency, for any smooth and compactly supported function $\kappa(\cdot)$ on $\mathbb{R}$ such that
(3.6) $\kappa(\nu)=1\quad\text{for}\quad\nu\in[-2\|H\|,2\|H\|]\,,$
we define the filtering function
(3.7) $\widehat{g}(\nu):=-\frac{i}{2}\tanh\left(-\frac{\beta\nu}{4}\right)\kappa(\nu)\,,$
and then have
(3.8) $G=\sum_{a\in\mathcal{A}}\sum_{\nu\in B_{H}}\hat{g}(\nu)\left(L_{a}^{\dagger}L_{a}\right)_{\nu}\,.$
Since $\widehat{g}(\nu)$ is an $L^{1}$-integrable function, its Fourier transform is well-defined:
(3.9) $g(t)=\frac{1}{2\pi}\int^{\infty}_{-\infty}\widehat{g}(\nu)e^{-it\nu}\,\mathrm{d}\nu\,,$
and it is direct to compute
(3.10) $G=\sum_{a\in\mathcal{A}}\sum_{\nu\in B_{H}}\widehat{g}(\nu)\left(L_{a}^{\dagger}L_{a}\right)_{\nu}=\int^{\infty}_{-\infty}g(t)H_{L}(t)\,\mathrm{d}t\,,$
where
(3.11) $H_{L}(t):=\sum_{a\in\mathcal{A}}e^{iHt}\left(L_{a}^{\dagger}L_{a}\right)e^{-iHt}\,.$
We emphasize that the choices of $q^{a}$ in (3.4) and $\kappa$ in (3.7) are key components in our algorithm and will be discussed in detail in Section 3.2.
With $G$ and $\{L_{a}\}_{a\in\mathcal{A}}$ be constructed in (3.10) and (3.3), respectively, the Lindbladian
@@
We have discussed the case where coupling operators $\{A^{a}\}_{a\in\mathcal{A}}$ are self-adjoint. In fact, one can generally consider a set of couplings such that $\{A^{a}\}_{a\in\mathcal{A}}=\{(A^{a})^{\dagger}\}_{a\in\mathcal{A}}$, and construct the corresponding jump operators $\{L_{a}\}_{a\in\mathcal{A}}$, the coherent term $G$, and the Lindbladian $\mathcal{L}$ as in Eqs. 3.3, 3.10 and 3.12, respectively. It is easy to see that the Lindblad dynamics defined in this way still satisfies the KMS detailed balance. Indeed, let $L_{a}$ and $L_{a,{\rm adj}}$ be the jumps associated with some $A^{a}$ and $(A^{a})^{\dagger}$ by (3.3) (without loss of generality, $A^{a}\neq(A^{a})^{\dagger}$). We then define self-adjoint operators
@@
3.2. Choice of weighting function $q(\nu)$ and $\kappa(\nu)$
In order to efficiently implement the jump operators $\{L_{a}\}_{a\in\mathcal{A}}$ in (3.3) and the coherent term $G$ in (3.10), we need to approximate the involved time–domain integrals by a numerical quadrature in a finite region. Specifically, we fix a truncation time $T>0$ and approximate the integral $\int^{\infty}_{-\infty}$ by $\int_{-T}^{T}$, which is then further approximated by numerical quadrature. To ensure efficiency, we need a small $T$, meaning that $f^{a}(t),g(t)$ should be smooth functions that decay rapidly as $\left\lvert t\right\rvert\to\infty$. To this end, we shall assume that $q^{a}$ and $\kappa$ are compactly supported Gevrey functions. We first recall the definition of Gevrey functions below [AHR17].
@@ 3.2. Choice of weighting function $q(\nu)$ and $\kappa(\nu)$
Some useful properties of Gevrey functions are collected in Appendix C. In particular, the product of two Gevrey functions is a Gevrey function (26); certain compositions of Gevrey functions are Gevrey functions (28); the Fourier transform of compactly supported Gevrey functions satisfies Paley-Wiener type estimates (29).
@@
For $\beta,S>0$, suppose that $q(\nu)$ is a weighting function of the form $q(\nu)=u(\beta\nu)w(\nu/S)$ with the following conditions:
@@
  • (Compact support) $w(\nu)=1$ when $|\nu|\leq 1/2$ and $w(\nu)=0$ when $|\nu|\geq 1$.
  • (Gevrey) There exists $\xi_{q},\xi_{u},\xi_{w}\geq 1$ and $s\geq 1$ such that
    $u(\nu)e^{-\nu/4}\in\mathcal{G}^{s}_{\xi_{q},\xi_{u}}(\mathbb{R})\,,\quad w\in\mathcal{G}^{s}_{\xi_{q},\xi_{w}}(\mathbb{R})\,.$
@@
Remark 14.
The above Gevrey functions with compact support are technical assumptions for our complexity analysis. Specifically, due to the properties of Gevrey functions discussed in Appendix C, we can demonstrate the rapid decay of $f^{a}$ and $g$ (see 30). Combining this with the compact support property of $\widehat{q}(\nu)$, we can apply the Poisson summation formula to control the quadrature error and normalization constant (see 15). In practical applications, we believe it is sufficient to choose a weighting function $q(\nu)$ with certain smoothness and decay to ensure efficient simulation of the quantum Gibbs sampler.
In the above assumption, $S$ is an adjustable parameter to control the support of $q$, and it is easy to see that $w(\nu/S)$ is supported on $[-S,S]$ and satisfies $w(\nu/S)\in\mathcal{G}^{s}_{\xi_{q},\xi_{w}/S}$. Then, for a weighting function $q$ in 13, there holds (26)
@@
(3.16) $q(\nu)e^{-\beta\nu/4}\in\mathcal{G}^{s,}_{\xi_{q}^{2},\beta\xi_{u}+\xi_{w}/S}(\mathbb{R})\,.$
Intuitively, the functions $u(\beta\nu)$ and $w(\nu/S)$ control the magnitude and support of $q$, respectively, which in turn controls the energy transition induced by jump operators $\{L_{a}\}_{a\in\mathcal{A}}$ (3.3). The existence of a bump function $w$ required in 13 is guaranteed by [AHR17, Corollary 2.8]: for any $s_{w}>1$, we can construct $w\in\mathcal{G}^{s_{w}}_{\xi_{q},\xi_{w}}$ with $\mathrm{supp}(w)\subset[-1,1]$ and $w(\nu)=1$ for $|\nu|\leq 1/2$.
We next discuss the choice of the weighting function $\kappa$ for the coherent term. Let $\{L_{a}\}_{a\in\mathcal{A}}$ be the jump operators defined via Eq. 3.3 with $q^{a}$ satisfying 13. We recall the representation (3.8) of $G$ and compute, for $a\in\mathcal{A}$,
$L_{a}^{\dagger}L_{a}$ $=\sum_{\nu,\nu^{\prime}\in B_{H}}e^{-\beta(\nu+\nu^{\prime})/4}q^{a}(\nu)\overline{q^{a}(\nu^{\prime})}\left(A^{a}_{\nu^{\prime}}\right)^{\dagger}A^{a}_{\nu}$
$=\sum_{\lambda_{j},\lambda_{j^{\prime}}\in{\rm Spec}(H)}\sum_{\lambda_{i}\in{\rm Spec}(H)}e^{-\beta(2\lambda_{i}-\lambda_{j}-\lambda_{j^{\prime}})/4}q^{a}(\lambda_{i}-\lambda_{j})\overline{q^{a}(\lambda_{i}-\lambda_{j^{\prime}})}\,P_{j^{\prime}}\left(A^{a}\right)^{\dagger}P_{i}A^{a}P_{j}\,,$
which, by (2.13), implies that for $\nu\in B_{H}$,
$(L_{a}^{\dagger}L_{a})_{\nu}$ $=\sum_{\lambda_{j^{\prime}}-\lambda_{j}=\nu}P_{j^{\prime}}\left(\sum_{\lambda_{i}\in{\rm Spec}(H)}e^{-\beta(2\lambda_{i}-\lambda_{j}-\lambda_{j^{\prime}})/4}q^{a}(\lambda_{i}-\lambda_{j})\overline{q^{a}(\lambda_{i}-\lambda_{j^{\prime}})}\,\left(A^{a}\right)^{\dagger}P_{i}A^{a}\right)P_{j}\,.$
By Eq. 3.15 from 13, we have that when $|\lambda_{j}-\lambda_{j}^{\prime}|>2S$, for any $\omega\in\mathbb{R}$,
$q^{a}(\omega-\lambda_{j})\overline{q^{a}(\omega-\lambda_{j^{\prime}})}=0\,.$
It follows that $(L_{a}^{\dagger}L_{a})_{\nu}=0$ if $|\nu|>2S$ and we can relax the constraint (3.6) to $\kappa(\nu)=1$ for $|v|\leq 2S$. Without loss of generality, we let $\kappa$ be
(3.17) $\kappa(\nu)=w(\nu/4S)\in\mathcal{G}^{s}_{\xi_{q},\xi_{w}/4S}\,,$
with $w(\cdot)$ given in 13.
We first prove that the filtering functions in the time domain $f^{a}(t)$ and $g(t)$ associated with $q^{a}$ and $\kappa$ given above decay rapidly (30). This further allows us to show that a simple quadrature scheme (trapezoidal rule) can efficiently approximate $\{L_{a}\}_{a\in\mathcal{A}}$ and $G$ with high accuracy, as summarized in 15. Specifically, given $M=2^{\mathfrak{m}-1}$ with $\mathfrak{m}\in\mathbb{N}_{+}$ and $\tau>0$, the quadrature points are given by
@@
The quadrature error can be controlled as follows, and its proof is given in Appendix D. For simplicity, in the main text, we assume that the parameters $\xi_{q}$, $\xi_{u}$, $\xi_{w}$ and $C_{1,u}$ for Gevrey functions in 13 are universal constants and include them in the asymptotic notations in related statements. We also assume $S\geq 1$ in the following discussion for simplicity. The explicit dependence on these parameters will be provided in Appendix D for completeness.
@@
Under 13, we assume $\beta>0$, $S\geq 1$, $\|A^{a}\|\leq 1$ for any $a\in\mathcal{A}$. When $\tau\leq\frac{\pi}{\left\lVert H\right\rVert+2S}$ and
(3.19) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(M-1)\tau=\Omega\left((\beta+2)\log(\beta+2)\right)\,,}$
it holds that
$\left\|L_{a}-\sum^{2M-1}_{m=0}f^{a}(t_{m})A^{a}(t_{m})\tau\right\|\leq C_{f}S\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2(\beta\xi_{u}+\xi_{w}/S)^{1/s}e}\right)\,,$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{f}=\mathcal{O}\big{(}(\beta+2)^{1/s}\big{)}\,,}$
@@
$\left\|G-\sum^{2M-1}_{m=0}g(t_{m})H_{L}(t_{m})\tau\right\|\leq C_{g}S|\mathcal{A}|\exp\left(-\frac{s((M-1)\tau)^{1/s}}{2e(\beta+\xi_{w}/4S)^{1/s}}\right)\,,$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{g}={\mathcal{O}}\big{(}(\beta+2)^{1/s}\log^{2}((\beta+2)S)\big{)}\,.}$
Here $\Omega(\cdot)$, $\mathcal{O}(\cdot)$ absorbs some constant depending on $s$, $\xi_{q}$, $\xi_{u}$, $\xi_{w}$ and $C_{1,u}$.
Thanks to 15, to approximately block encode $\{L_{a}\}$ and $G$, it suffices to construct block-encodings for the discretized quantities
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum^{2M-1}_{n=0}g(t_{n})e^{iHt_{n}}\left(\sum_{a\in\mathcal{A}}L^{\dagger}_{a}L_{a}\right)e^{-iHt_{n}}\tau\,.}$
In our algorithm, we construct these two block encodings using LCU (see Appendix B). This utilizes block encodings of $A^{a}$ (3.24), controlled Hamiltonian simulation (3.25), and prepare oracles for $f$ (Eqs. 3.27 and 3.28) or $g$ (Eqs. 3.30 and 3.31). The detailed constructions are presented in the next subsection (see (3.33) and (3.34)).
@@
Bounding the approximation error for $\{L_{a}\}_{a\in\mathcal{A}}$ and $G$ in the operator norm is a nontrivial task. [CKBG23] introduces a “rounding Hamiltonian” technique to bound the quadrature error in the frequency domain. By choosing weighting functions in 13, we can use the Poisson summation formula to simplify the quadrature error analysis. In particular, we can bound the quadrature error in the time domain without using the “rounding Hamiltonian” technique.
Now we provide two specific examples of $q$ satisfying 13.
@@
    where $u(\nu)e^{-\frac{\nu}{4}}=e^{-\frac{\sqrt{1+\nu^{2}}+\nu}{4}}\in\mathcal{G}^{1}_{1,7/2}$ is a Gevrey function of order $1$ and its derivative is $L^{1}$-integrable; see 28. When $\beta\gg 1$, $q(\nu)$ in Eq. 3.20 gives a smoothed version of the Metropolis-type filter (similar to the Glauber-type filter):